OpenStudy (anonymous):

Am i right? One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there? Let's assume that HM = Human and HR = Horse HM + HR = 74 2HM + 4HR = 196 (2HM + 4HR) - (2 HM + 2HR) = 196 - 148 2HR = 48 HR = 24 HM + (24) = 74 HM = 74 - 24 HM = 50 24 horses and 50 humans

OpenStudy (jdoe0001):

(2HM + 4HR) - (2 HM + 2HR) = 196 - 148 <--- same - same = 0

OpenStudy (jdoe0001):

ohh ... my bad... they're not the same ..hmmm ok

OpenStudy (jdoe0001):

anyhow, yes you're correct

OpenStudy (anonymous):

k thanks!!!!!:)

OpenStudy (jdoe0001):

though if it was me, I could tell a horse head from a human head, I'd have just counted them separately =)

OpenStudy (anonymous):

lol I would too